# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multimodal Large Language Models

### v7

@@ 1. Introduction
Deep generative models, such as Variational Autoencoders (VAEs) (Kingma and Welling, 2013) and diffusion models (Rombach et al., 2022), have become a state-of-the-art approach in various generation tasks (Ho et al., 2022; Yang et al., 2023). These methods effectively leverage latent variables to learn underlying data distributions and generate high-quality samples by capturing the underlying structure of high-dimensional data in a low-dimensional semantic space. As the latent variables represent all the information in a lower dimension, they can be considered as an effective abstraction of key factors in the data. Therefore, it is critical to develop methods for automatically decomposing and explaining meaningful latent dimension semantics given a pretrained generative model and its inherent inductive biases, as illustrated in Figure 1. Inductive biases are often enforced over latent variables in deep generative models. For instance, disentanglement is a rule of thumb which enforces orthogonality among different latent variables (Ding et al., 2020). Moreover, sometimes latent variables can be grouped, leading to combination bias (Klys et al., 2018). More recently, the desire for controllability in deep generative models, where latent variables are associated with specific properties of interest (Wang et al., 2024), has given rise to conditional bias. Incorporating inductive biases aligned with the actual facts can reduce the hallucination in explaining the latent variables in deep generative models (Wu et al., 2024).
@@
The field of explainable artificial intelligence (XAI) has extensively investigated the interpretation of machine learning models (Adadi and Berrada, 2018; Zhu et al., 2021). However, interpreting latent variables in deep generative models remains underexplored. Machine learning model explanation, a.k.a., post-hoc explanation, can be categorized into global and local explanations (Gao et al., 2024). Global explanations focus on elucidating the entire model, while local explanations target the reasoning behind specific predictions. Global explanations are more challenging, with existing work mostly emphasizing attributions to identify which features are most important for model decision-making (Saleem et al., 2022). The missing piece is understanding the meaning of features when they are unknown, which is very common in deep generative models. More recently, one category of global explanation methods, called concept-based explanations, aims to generate more human-understandable concepts as explanations (Poeta et al., 2023). However, current concept-based methods often rely on human heuristics or predefined concept and feature space, limiting the expressiveness of the explanations and falling short of achieving truly automatic explanation generation (Koh et al., 2020; Bai et al., 2022).
@@ 2. Related Work
Deep Generative Models. Deep generative models are essential for modeling complex data distributions. Variational Autoencoders (VAEs) are prominent in this area, introduced by Kingma and Welling (Kingma and Welling, 2013). VAEs encode input data into a latent space and decode it back, optimizing a balance between reconstruction error and the Kullback-Leibler divergence (Rezende et al., 2014). They have diverse applications, including image generation (Yan et al., 2016), and anomaly detection (An and Cho, 2015).
@@
Diffusion models, proposed by Ho et al. (Ho et al., 2020), generate data by a diffusion process that gradually adds noise to the data and then learns to reverse this process to recover the original data. These models have achieved high-fidelity image generation, surpassing generative adversarial networks (GANs) in quality and diversity. Latent diffusion models allow the model to operate in a lower-dimensional space, which significantly reduces computational requirements while maintaining the quality of the generated samples (Rombach et al., 2022). Advances have made them applicable to text-to-image synthesis (Nichol and Dhariwal, 2021), and audio generation (Kong et al., 2020).
@@
Latent Variable Manipulation and Explanations. Manipulating latent variables in generative models like VAEs and diffusion models is an important technique for editing and enhancing generated images. A key method is latent traverse, which involves traversing different values of latent variables to achieve diverse manipulations in the generated outputs. This technique allows for precise control over the attributes in generated images, enabling adjustments (Chen et al., 2016). For example, latent traverse has been effectively employed to disentangle and control various attributes in generated images (Brock et al., 2016; Zhu et al., 2016). However, latent traverse is often used for visualization and editing purposes. It has not yet been widely explored as a tool for explaining the underlying latent space. Their explanations primarily rely on using predefined training attributes as text labels or manually adding explanations (Shen et al., 2020; Esser et al., 2020). Some concept-based models control latent variables in generative models using category concepts to generate data (Tran et al., 2022; Bouchacourt et al., 2018). However, these approaches cannot automatically generate free-form textual explanations.
@@
More recently, Multimodal Large Language Models (MLLMs) integrate diverse data modalities, enhancing their ability to understand and generate complex information (Yin et al., 2023; Bai et al., 2024). Notable models include GPT-4o, which extends GPT-4v with better visual capabilities (OpenAI, 2024), and Gemini that is a family of highly capable multimodal models (Team et al., 2024). We envision leveraging MLLMs to automatically generate explanations for latent variables and incorporate their inductive biases to reduce hallucination. Our work focuses on: (1) how to decompose the inductive bias formulas to automate the manipulation of latent variables, (2) how to develop prompts that are aligned with the underlying inductive biases of generative models and can be easily understood by MLLMs, and (3) how to evaluate the quality of the generated explanations for latent variables.
@@ 3. Preliminaries and Problem Formulation
Deep Generative Models. These are a class of models that learn a mapping between observations $\mathbf{x}$ and underlying latent factors $\mathbf{z}$. These models are widely used in deep generative models such as VAEs and latent diffusion models. VAEs, for instance, introduce a probabilistic approach to encoding data by maximizing the evidence lower bound (ELBO) (Kingma and Welling, 2013):
@@
Inductive Bias in Latent Variables. Inductive biases are usually imposed on latent variables to enhance the performance and interpretability of deep generative models. These inductive biases in deep generative models can be categorized into three common types: Disentanglement Bias: It enforces orthogonality among different latent variables, ensuring that each latent variable captures a distinct factor of variation in the data (Ding et al., 2020). Combination bias: Sometimes latent variables are grouped, leading to biases in how they interact and combine to represent complex data structures (Klys et al., 2018). Conditional Bias: It emphasizes the relationship between specific properties of interest and the corresponding latent variables (Wang et al., 2024).
@@ 4.2.2. From Disentanglement Bias to Prompts
Disentanglement bias refers to the model’s ability to separate independent factors in the data (Ding et al., 2020; Wu et al., 2023). The formula representing this bias focuses on ensuring that different latent variables correspond to different independent underlying factors. Independent factors would be invariant with respect to one another (Ridgeway, 2016). By disentangling these factors, researchers can better understand the underlying structure of the data and improve the model’s performance on tasks such as representation learning.
@@ 4.2.3. From Combination Bias to Prompts
Combination bias involves understanding how different latent variables interact within groups and remain independent across groups (Klys et al., 2018). This bias is significant as it helps in identifying how combinations of factors contribute to the overall data generation process. Recognizing these interactions enables researchers to design models that can generate more complex and realistic data by capturing intricate relationships within the data.
@@ 4.2.4. From Conditional Bias to Prompts
Conditional bias focuses on the relationship between specific properties of interest and the corresponding latent variables (Wang et al., 2024). This bias is important because it allows models to generate data conditioned on particular attributes, enhancing the model’s ability to produce targeted and controlled outputs.
@@ 4.3. Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ within these three bias using in-context learning.
@@ 4.5. Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses (Lin et al., 2023) and image explanations (Zhao et al., 2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses $\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. Our final explanation $\hat{r}$ is the response that has the highest pairwise cosine similarity with other responses if the latent variable is interpretable. Otherwise, $\hat{r}$ will be “No clear explanation”. The implementation details are in Section 5.5.
@@ 5.1. Dataset
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ (Karras et al., 2018), AFHQ (Choi et al., 2020), LSUN-Church (Yu et al., 2015) for the unconditional and conditional diffusion models, and CelebA-HQ (Karras et al., 2018), 3DShapes (Burgess and Kim, 2018), and dSprites (Matthey et al., 2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
@@ 5.2. Models and Baselines
Our evaluation benchmarks our proposed LatentExplainer framework against three state-of-the-art multimodal models with strong vision-language reasoning capabilities: GPT-4o (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024). We employ GPT-4o, Gemini 1.5 Pro, and Claude 3.5 as a zero-shot baseline, comparing it with the addition of our Latent Exlainer with both the inductive bias prompt and uncertainty quantification included.
@@ 5.3. Generative Models under Inductive Biases
We explore the latent space in generative models that satisfy the aforementioned three types of inductive biases. For each type, we present the relevant generative models that align with the corresponding inductive bias: (1) Disentanglement Bias: $\beta$-TCVAE (Chen et al., 2018) explicitly penalizes the total correlation of the latent variables to disentangle the latent variables. Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) adds Gaussian noise independently at each timestep in the forward process and eventually transforms into pure Gaussian noise, in which the covariance matrix is diagonal. This assumes the latent factors are independent; (2) Combination Bias: CSVAE (Klys et al., 2018) has two groups of latent variables $z$ and $w$, where $z$ and $w$ are uncorrelated and the latent variables within the group are correlated; (3) Conditional Bias: CSVAE also satisfies conditional bias because one group of latent variables $w$ is associated with the properties while the other group of latent variables $z$ minimizes the mutual information with the properties. Stable Diffusion (Rombach et al., 2022) is a latent diffusion model to generate images conditioned on prompts.
@@ 5.5. Uncertainty-aware Implementations
We follow (Zhu et al., 2024) to denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
@@ 5.7. Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), SPICE (Anderson et al., 2016), BERTScore(Zhang et al., 2019), and BARTScore (Yuan et al., 2021) as the automated metrics to assess the generated explanations. BLEU, and ROUGE-L are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.

### v8

@@ 1. Introduction
Deep generative models, such as Variational Autoencoders (VAEs) (kingma2013auto) and diffusion models (rombach2022high), have become a state-of-the-art approach in various generation tasks (ho2022cascaded; yang2023diffusion). These methods effectively leverage latent variables to learn underlying data distributions and generate high-quality samples by capturing the underlying structure of high-dimensional data in a low-dimensional semantic space. As the latent variables represent all the information in a lower dimension, they can be considered as an effective abstraction of key factors in the data. Therefore, it is critical to develop methods for automatically decomposing and explaining meaningful latent dimension semantics given a pretrained generative model and its inherent inductive biases, as illustrated in Figure 1. Inductive biases are often enforced over latent variables in deep generative models. For instance, disentanglement is a rule of thumb which enforces orthogonality among different latent variables (ding2020guided). Moreover, sometimes latent variables can be grouped, leading to combination bias (klys2018learning). More recently, the desire for controllability in deep generative models, where latent variables are associated with specific properties of interest (wang2024controllable), has given rise to conditional bias. Incorporating inductive biases aligned with the actual facts can reduce the hallucination in explaining the latent variables in deep generative models (wu2024no).
@@
The field of explainable artificial intelligence (XAI) has extensively investigated the interpretation of machine learning models (adadi2018peeking; zhu2021development). However, interpreting latent variables in deep generative models remains underexplored. Machine learning model explanation, a.k.a., post-hoc explanation, can be categorized into global and local explanations (gao2024going). Global explanations focus on elucidating the entire model, while local explanations target the reasoning behind specific predictions. Global explanations are more challenging, with existing work mostly emphasizing attributions to identify which features are most important for model decision-making (saleem2022explaining). The missing piece is understanding the meaning of features when they are unknown, which is very common in deep generative models. More recently, one category of global explanation methods, called concept-based explanations, aims to generate more human-understandable concepts as explanations (poeta2023concept). However, current concept-based methods often rely on human heuristics or predefined concept and feature space, limiting the expressiveness of the explanations and falling short of achieving truly automatic explanation generation (koh2020concept; bai2022concept).
@@ 2. Related Work
Deep Generative Models. Deep generative models are essential for modeling complex data distributions. Variational Autoencoders (VAEs) are prominent in this area, introduced by Kingma and Welling (kingma2013auto). VAEs encode input data into a latent space and decode it back, optimizing a balance between reconstruction error and the Kullback-Leibler divergence (rezende2014stochastic). They have diverse applications, including image generation (yan2016attribute2image), and anomaly detection (an2015variational).
@@
Diffusion models, proposed by Ho et al. (ho2020denoising), generate data by a diffusion process that gradually adds noise to the data and then learns to reverse this process to recover the original data. These models have achieved high-fidelity image generation, surpassing generative adversarial networks (GANs) in quality and diversity. Latent diffusion models allow the model to operate in a lower-dimensional space, which significantly reduces computational requirements while maintaining the quality of the generated samples (rombach2022high). Advances have made them applicable to text-to-image synthesis (nichol2021glide), and audio generation (kong2020diffwave).
@@
Latent Variable Manipulation and Explanations. Manipulating latent variables in generative models like VAEs and diffusion models is an important technique for editing and enhancing generated images. A key method is latent traverse, which involves traversing different values of latent variables to achieve diverse manipulations in the generated outputs. This technique allows for precise control over the attributes in generated images, enabling adjustments (chen2016infogan). For example, latent traverse has been effectively employed to disentangle and control various attributes in generated images (brock2016neural; zhu2016generative). However, latent traverse is often used for visualization and editing purposes. It has not yet been widely explored as a tool for explaining the underlying latent space. Their explanations primarily rely on using predefined training attributes as text labels or manually adding explanations (shen2020interpreting; esser2020disentangling). Some concept-based models control latent variables in generative models using category concepts to generate data (tran2022unsupervised; bouchacourt2018multi). However, these approaches cannot automatically generate free-form textual explanations.
@@
More recently, Multimodal Large Language Models (MLLMs) integrate diverse data modalities, enhancing their ability to understand and generate complex information (yin2023survey; bai2024beyond). Notable models include GPT-4o, which extends GPT-4v with better visual capabilities (openai2024gpt4o), and Gemini that is a family of highly capable multimodal models (team2024gemini). We envision leveraging MLLMs to automatically generate explanations for latent variables and incorporate their inductive biases to reduce hallucination. Our work focuses on: (1) how to decompose the inductive bias formulas to automate the manipulation of latent variables, (2) how to develop prompts that are aligned with the underlying inductive biases of generative models and can be easily understood by MLLMs, and (3) how to evaluate the quality of the generated explanations for latent variables.
@@ 3. Preliminaries and Problem Formulation
Deep Generative Models. These are a class of models that learn a mapping between observations $\mathbf{x}$ and underlying latent factors $\mathbf{z}$. These models are widely used in deep generative models such as VAEs and latent diffusion models. VAEs, for instance, introduce a probabilistic approach to encoding data by maximizing the evidence lower bound (ELBO) (kingma2013auto):
@@
Inductive Bias in Latent Variables. Inductive biases are usually imposed on latent variables to enhance the performance and interpretability of deep generative models. These inductive biases in deep generative models can be categorized into three common types: Disentanglement Bias: It enforces orthogonality among different latent variables, ensuring that each latent variable captures a distinct factor of variation in the data (ding2020guided). Combination bias: Sometimes latent variables are grouped, leading to biases in how they interact and combine to represent complex data structures (klys2018learning). Conditional Bias: It emphasizes the relationship between specific properties of interest and the corresponding latent variables (wang2024controllable).
@@ 4.2.2. From Disentanglement Bias to Prompts
Disentanglement bias refers to the model’s ability to separate independent factors in the data (ding2020guided; wu2023uncovering). The formula representing this bias focuses on ensuring that different latent variables correspond to different independent underlying factors. Independent factors would be invariant with respect to one another (ridgeway2016survey). By disentangling these factors, researchers can better understand the underlying structure of the data and improve the model’s performance on tasks such as representation learning.
@@ 4.2.3. From Combination Bias to Prompts
Combination bias involves understanding how different latent variables interact within groups and remain independent across groups (klys2018learning). This bias is significant as it helps in identifying how combinations of factors contribute to the overall data generation process. Recognizing these interactions enables researchers to design models that can generate more complex and realistic data by capturing intricate relationships within the data.
@@ 4.2.4. From Conditional Bias to Prompts
Conditional bias focuses on the relationship between specific properties of interest and the corresponding latent variables (wang2024controllable). This bias is important because it allows models to generate data conditioned on particular attributes, enhancing the model’s ability to produce targeted and controlled outputs.
@@ 4.3. Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ within these three biases using in-context learning.
@@ 4.5. Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses (lin2023generating) and image explanations (zhao2024due). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses $\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. Our final explanation $\hat{r}$ is the response that has the highest pairwise cosine similarity with other responses if the latent variable is interpretable. Otherwise, $\hat{r}$ will be “No clear explanation”. The implementation details are in Section 5.5.
@@ 5.1. Dataset
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ (karras2018progressive), AFHQ (choi2020stargan), LSUN-Church (yu15lsun) for the unconditional and conditional diffusion models, and CelebA-HQ (karras2018progressive), 3DShapes (3dshapes18), and dSprites (dsprites17) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
@@ 5.2. Models and Baselines
Our evaluation benchmarks our proposed LatentExplainer framework against three state-of-the-art multimodal models with strong vision-language reasoning capabilities: GPT-4o (openai2024gpt4o), Gemini 1.5 Pro (team2024gemini), and Claude 3.5 Sonnet (claude35sonnet2024). We employ GPT-4o, Gemini 1.5 Pro, and Claude 3.5 as a zero-shot baseline, comparing it with the addition of our Latent Exlainer with both the inductive bias prompt and uncertainty quantification included.
@@ 5.3. Generative Models under Inductive Biases
We explore the latent space in generative models that satisfy the aforementioned three types of inductive biases. For each type, we present the relevant generative models that align with the corresponding inductive bias: (1) Disentanglement Bias: $\beta$-TCVAE (chen2018isolating) explicitly penalizes the total correlation of the latent variables to disentangle the latent variables. Denoising Diffusion Probabilistic Model (DDPM) (ho2020denoising) adds Gaussian noise independently at each timestep in the forward process and eventually transforms into pure Gaussian noise, in which the covariance matrix is diagonal. This assumes the latent factors are independent; (2) Combination Bias: CSVAE (klys2018learning) has two groups of latent variables $z$ and $w$, where $z$ and $w$ are uncorrelated and the latent variables within the group are correlated; (3) Conditional Bias: CSVAE also satisfies conditional bias because one group of latent variables $w$ is associated with the properties while the other group of latent variables $z$ minimizes the mutual information with the properties. Stable Diffusion (rombach2022high) is a latent diffusion model to generate images conditioned on prompts.
@@ 5.5. Uncertainty-aware Implementations
We follow (zhu2024explaining) to denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
@@ 5.7. Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU (papineni2002bleu), ROUGE-L (lin-2004-rouge), SPICE (anderson2016spice), BERTScore(zhang2019bertscore), and BARTScore (yuan2021bartscore) as the automated metrics to assess the generated explanations. BLEU, and ROUGE-L are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.